How fast is the surface area of a spherical balloon increasing when the radius is 10 cm and

the volume is increasing at 15 cm3/sec? [7]

2. Let f be the function defined by

f (x) = xe-x:

(a) Determine the y–intercept.

(b) Determine the horizontal and vertical asymptotes.

(c) Use the sign pattern for f' (x) to determine

(i) the interval(s) over which f rises and where it falls;

(ii) the local extrema.

(d) Use the sign pattern for f'' (x) to determine

(i) where the graph of f is concave up and where it is concave down;

(ii) the inflection point(s) (if any).


1
Expert's answer
2021-05-07T09:41:03-0400

1.

Let r, v and s be the radius, the volume and the surface area of a sphere at any instant t respectively.


Volume of sphere="v=\\frac{4}{3}\u200b\u03c0r^3"

Then,

"v=\\frac{4}{3}\u200b\u03c0r^3\\newline\n\n\\frac{dv}{dt}=4\u200b\u03c0r^2.\\frac{dr}{dt}\u200b\\newline\n\n\\text{given} \\frac{dv}{dt}\u200b=15 \\text { cm}^3 \\text{\/sec and} r=10 \\text{cm}\\newline\n\n\\text{Therefore}, 15=4\u200b\u03c0\u00d7(10)^2\u00d7\\frac{dr}{dt}\u200b\\newline\n\n\\frac{dr}{dt}=\\frac{15}{400\u03c0}=\\frac{3}{80\u03c0}\\newline"

Now, surface area="s=4\u03c0r^2"

Then,

"s=4\u03c0r^2\\newline\n\\frac{ds}{dt}\u200b=8\u03c0r\\frac{dr}{dt}\u200b\\newline\n\nor, \\frac{ds}{dt}\u200b=8\u00d7\u03c0\u00d710\u00d7\\frac{3}{80\u03c0}\u200b \\hspace{1cm} [as\\space r=10, \\frac{dr}{dt}\u200b=\\frac{3}{80\u03c0}\u200b]\\newline\n\n \\frac{ds}{dt}\u200b=3 \\text{cm}^2\\text{\/sec}.\\newline\n\\text{Thus, the rate change of surface area 3cm}^2\\text{\/sec.}"


2.

Given, "f(x)=xe^{-x}" .


(a)

"put \space x=0,\newline
y(0)=0"

Thus, the y-intercept is y=0.


(b)

For horizontal asymptotes,

"lim_{x \to ∞} xe^{-x}=lim_{x \to ∞} \frac{x}{e^{x}}\hspace{1cm} ( intermidiate form)\newline
\text{By L'Hospital rule,}\newline
lim_{x \to ∞} \frac{1}{e^{x}}=0"

Thus, horizontal asymptotes is y=0.

For vertical asymptotes,

When y "\to" ∞.

Therefore, "f(x)=xe^{-x}=\frac{x}{e^x}"

When "e^{x} =0" then y "\to" to ∞. But "e^{x} \neq 0, \forall x".

Therefore, there is no vertical asymptote.


(c)

"f'(x)=e^{-x}(1-x)\newline
\text{Now, f'(x)=0 at x=1.}\newline
\text{Therefore, intervals are} (-∞,1) \text{and} (1,∞).\newline
(i)\newline
In (-∞,1), f'(x)<0 \implies \text{f rises} \newline
\text{and in} (1,∞) f'(x)>0 \implies \text{ f fall }.\newline
(ii)\newline
\text{Since, there is only one extreme point. And, it is global maxima.}"


(d)

"f''(x)=e^{-x}(x-2)\newline
\text{Now, f''(x)=0 at x=2.}\newline
\text{Therefore, intervals are} (-∞,2) \text{and} (2,∞).\newline
(i)\newline
In (-∞,2), f''(x)<0 \implies \text{f is concave down} \newline
\text{and in} (2,∞) f''(x)>0 \implies \text{ f is concave up}.\newline
(ii)\newline
\text{Infection point is x=2,because there f''=0.} \newline \text{And, f changes from concave down to concave up.}"